a high school graduating class is made up of 455 students. there are 67 more girls than boys. how many boys and girls are in the class?
Answer by Cromlix(4381)

Hi there,
Make the number of boys = 'x'
Girls = x + 67
Therefore x + (x + 67) = 455
2x + 67 = 455
2x = 455 - 67
2x = 388
x = 194
Boys = 194
Girls = 261
Hope this helps:-)